A compact car has a maximum acceleration of 3.0m/s2 when it carries only the driver and has a total mass of 1100kg.

What is its maximum acceleration after picking up four passengers and their luggage, adding an additional 400 kg of mass?

Express your answer to two significant figures and include the appropriate units.
Answer #1

Mass of Car = 1100 kg, a = 3.0 m/s2

Force = m * a = 1100 * 3.0 = 3300N

Weight =1100 + 400 = 1500 kg

Force = Mass * acceleration

3300 = 1500 * a

a = 2.2 m/s2